# Cross-talk in superconducting qubit lattices with tunable couplers – comparing transmon and fluxonium architectures

### v1

@@ III.3 C-shunt flux couplers
$\hat{H}_{\text{CSF}}$ $\approx\omega\,\hat{b}^{\dagger}\hat{b}+\frac{\alpha}{2}\hat{b}^{\dagger}\hat{b}^{\dagger}\hat{b}\hat{b}$
$\hskip 14.22636pt+3K_{3}\left(\frac{E_{C}}{K_{2}}\right)^{\mathclap{\frac{3}{2}}}(\hat{b}^{\dagger}+\hat{b}+\hat{b}^{\dagger}\hat{b}^{\dagger}\hat{b}+\hat{b}^{\dagger}\hat{b}\hat{b})\,,$ (12)
@@
Previous work has demonstrated that the nearest-neighbor $ZZ$ coupling can be fully removed even when spectator qubits are included, although the ideal parameter points may change noticeably compared with isolated qubit pairs [12]. In order to have a fair comparison with transmon couplers, we first pick the parameters based on isolated qubit dimers as before. We again set $E_{C}=0.15$ GHz for the couplers and keep the flux parameter fixed at $\phi_{\text{ext}}=0.51$, which is slightly detuned from the point of maximum anharmonicity $\alpha$ and minimum frequency $\omega$ to allow for a fine tuning of $\omega$ in both directions at a later stage. The condition of vanishing $ZZ$ coupling defines a curve in the $\gamma$-$E_{J}$ plane. We minimize the parameter
@@
Figure. 5 shows the interaction coefficients $w_{n}$ for the qubit ladder with model parameters determined in this manner. As expected, the overall reduction of $ZZ$ interactions is significantly stronger than for tunable transmon couplers, with the nearest-neighbor terms remaining below 202020\,20kHz for $T\leq 20\,$MHz. This corresponds to a reduction of at least one order of magnitude compared to conventional tunable couplers. The specific long-ranged terms discussed for the models with static coupling and tunable transmon couplers are also strongly suppressed, which is perhaps related to an enhanced localization of the dressed qubits due to the minimization of the delocalization parameter $\epsilon$ in Eq. (20). However, the IPR does not differ much compared to the model with transmon couplers (Fig. 4), likely because our parameter choice results in very similar hopping amplitudes for the bosons in both models, and is in fact slighly lower for $T\lesssim 12$ MHz. The mutual information, on the other hand, is noticeably smaller, especially at larger $T$. Together with the results for the interaction coefficients $w_{n}$, this suggests that the C-shunt flux couplers are significantly more effective in reducing the buildup of longer-ranged correlations with $T$ than transmon couplers (whose parameters are optimized to minimize the $ZZ$ interaction). To fully confirm that this is a general difference between the two coupling schemes, one would have to check additional frequency patterns, however.
Figure 6: (a) Approximate $ZZ$ coupling $\zeta$ defined in Eq. (21) between qubits $3$ and $4$ as a function of external flux and capacitive coupling strength. (b) $ZZ$ coupling for the isolated qubit pair. The avoided level crossing responsible for the jump in $\zeta$ is shown in panel (c).
@@
where $E_{n_{1}n_{2}}^{\prime}$ is the energy of the state, in which the qubits connected to the coupler are in the states $n_{1}$ and $n_{2}$, and all other qubits are in the 00 state. This becomes equal to the absolute values of nearest-neighbor couplings in Eq. (2) when higher-order terms involving 3 or more qubits vanish. As an example, Fig. 6(a) displays $\zeta$ as a function of $\phi_{\text{ext}}$ and $T$ for the qubit pair $2$ and $3$. For small $T$, the difference compared to the isolated qubit dimer [Fig. 6(b)] is mostly a change of the optimal value for $\phi_{\text{ext}}$. Around $T\approx 10$ MHz, however, there is a sudden increase of $\zeta$ caused by an avoided crossing between the eigenstates for $n=01100000$ and $n=00110000$ [Fig. 6(c)], showing that the effect of spectator qubits cannot always be accounted for by a simple shift of $\phi_{\text{ext}}$. It should also be noted that the optimization only addresses the nearest-neighbor $ZZ$ interactions, so we can not reasonably expect it to remove all stray couplings for larger $T$, where longer-ranged and multi-qubit terms become important. Nevertheless, the DMRG simulations confirmed that minimizing $\zeta$ independently for each coupler while keeping the remaining model parameters fixed is sufficient to bring all interaction strengths $|w_{n}|$ below 50 Hz for weak capacitive coupling $T\leq 8$ MHz.
@@ IV Fluxonium with transmon couplers
The qubit parameters need to be chosen carefully to ensure that the envisioned two-qubit gates are possible while keeping the $ZZ$-crosstalk low. For long coherence times, $E_{L}$ should also not be too large. Based on these considerations, we propose the set of parameters given in Table 2. The strengths of the capacitive couplings are again uniform. For each $T$, we first determine the value of $T_{c}$ that minimizes the average $ZZ$ coupling between nearest-neighbor qubits when treating the qubit dimers as isolated (see Fig. 7). We then use these parameteres for the MPS simulation of the full system.
Figure 8 shows the obtained nearest-neighbor $ZZ$ couplings and the average IPR. The $ZZ$ couplings are on the order of a few kHz and deviate only slightly from the results for the isolated qubit dimers. Moreover, the additional longer-ranged interactions and those involving three or more qubits have coefficients $w_{n}$ with magnitude below 10 Hz, and are thus negligible. The IPR remains at a large value $>0.9$ for all considered parameters, indicating that the computational basis states remain strongly localized. This can likely be attributed to the fact that the capacitive coupling only weakly mixes the $|0\rangle$ and $|1\rangle$ states of the fluxonium, and that the additional transmons have a much larger transition frequency, so that the hybridization with the couplers is also suppressed.
@@
Taking the corresponding eigenstates of the full system into account one can define an effective Hamiltonian similar to (2) by including terms with operators $\hat{P}_{i}=\sum_{n\in\{0,1,2\}^{N}}\delta_{n_{i}2}|\tilde{n}\rangle\langle\tilde{n}|$. As an example, Fig. 8(c) shows the coefficients for interactions of the form ${\propto}\hat{Z}_{i}\hat{P}_{j}$, which would be relevant when implementing the CZ gate mentioned above. The calculation only requires eigenstates with at most one qubit in the $|2\rangle$ state, significantly reducing the number of DMRG simulations. For our model, we find that the nearest-neighbor couplings are on the order of 111\,1 MHz, while the next-nearest neighbor ones reach values around a few tens of kHz. These stronger interactions are expected because of the reduced localization of the states and the fact that the parameters were chosen to minimize the $ZZ$ coupling specifically.
Figure 7: $ZZ$ coupling averaged over decoupled subsystems of nearest-neighbor fluxonium qubits and the connecting transmons as a function of the coupling strengths $T$ and $T_{c}$. The line shows the $T$-$T_{c}$ curve used for the MPS simulations of the full system.
@@
V Conclusion
Building a superconducting quantum computer requires achieving fast gate operations while minimizing unwanted crosstalk across the qubit chip. Although increasing coupling strengths enables faster gates, it also amplifies crosstalk and may even induce chaotic instabilities in larger systems, making it challenging to strike an optimal balance. We therefore used density-matrix renormalization-group simulations to investigate how well different schemes for superconducting qubit devices are able to suppress the static $ZZ$ interaction when the system is scaled up.
In agreement with previous works, we found that lattices of fixed-frequency transmons are limited to small capacitive couplings on the order of a few MHz, before the nearest-neighbor $ZZ$ interaction becomes prohibitively large. Although the addition of tunable couplers can reduce the effective interactions between qubits significantly, the qubits will still delocalize at sufficiently large capacitive couplings, and we indeed observe emerging long-range interactions due to near-resonant next-nearest neighbor qubits. Avoiding such frequency collisions should therefore also be beneficial for architectures with tunable couplers.
@@ V Conclusion
For fluxonium qubits, we showed that a transmon-based tunable coupler can suppress $ZZ$ crosstalk to the level of $\sim 1\,$kHz. However, to contextualize this result, particularly in comparison to transmon architectures, we must assess the coupling strengths required for implementing gates with similar speeds. In Ref. [14], a coupling strength of $T\approx 120\,$MHz was used to implement a 60 ns CZ gate for two Fluxoniums. In contrast, a similar CZ gate in a transmon system with transmon tunable couplers, as demonstrated in Ref. [9], was achieved in a 38 ns gate time with $T\approx 20\,$MHz. Notice that both gates were studied in isolated two-qubit systems. Our simulations suggest that in a scaled-up qubit layout, residual crosstalk in a fluxonium system would be an order of magnitude lower, indicating that crosstalk might not be the primary bottleneck for scaling fluxonium-based architectures. It should be noted, however, that by treating the nearest-neighbor coupling as an independent parameter, we did not account for capacitive loading [38], which can become a problem when scaling up the number of qubits as it limits the attainable coupling strengths for fluxoniums with higher connectivity. Similarly, we neglected the longer-ranged terms in the Hamiltonian that should become more relevant in both transmon and fluxonium systems as the coupling capacitances increase. In principle, such terms could be easily included in the DMRG simulations, and quantifying their effect on the localization properties might be an interesting direction for future study. It would also be possible to extend the MPS approach used here to time-dependent simulations [39], e.g., to investigate the effect of stray couplings during gate operations [40].

### v2

@@ III.3 C-shunt flux couplers
$\hat{H}_{\text{CSF}}$ $\approx\nu\,\hat{b}^{\dagger}\hat{b}+\frac{\alpha}{2}\hat{b}^{\dagger}\hat{b}^{\dagger}\hat{b}\hat{b}$
[…]
@@
Previous work has demonstrated that the nearest-neighbor $ZZ$ coupling can be fully removed even when spectator qubits are included, although the ideal parameter points may change noticeably compared with isolated qubit pairs [12]. In order to have a fair comparison with transmon couplers, we first pick the parameters based on isolated qubit dimers as before. We again set $E_{C}=0.15$ GHz for the couplers and keep the flux parameter fixed at $\phi_{\text{ext}}=0.51$, which is slightly detuned from the point of maximum anharmonicity $\alpha$ and minimum frequency $\nu$ to allow for a fine tuning of $\nu$ in both directions at a later stage. The condition of vanishing $ZZ$ coupling defines a curve in the $\gamma$-$E_{J}$ plane. We minimize the parameter
@@
Figure. 5 shows the interaction coefficients $w_{\bm{n}}$ for the qubit ladder with model parameters determined in this manner. As expected, the overall reduction of $ZZ$ interactions is significantly stronger than for tunable transmon couplers, with the nearest-neighbor terms remaining below 2020\,kHz for $T\leq 20\,$MHz. This corresponds to a reduction of at least one order of magnitude compared to conventional tunable couplers. The specific long-ranged terms discussed for the models with static coupling and tunable transmon couplers are also strongly suppressed, which is perhaps related to an enhanced localization of the dressed qubits due to the minimization of the delocalization parameter $\epsilon$ in Eq. (20). However, the IPR does not differ much compared to the model with transmon couplers (Fig. 4), likely because our parameter choice results in very similar hopping amplitudes for the bosons in both models, and is in fact slightly lower for $T\lesssim 12$ MHz. The mutual information, on the other hand, is noticeably smaller, especially at larger $T$. Together with the results for the interaction coefficients $w_{\bm{n}}$, this suggests that the C-shunt flux couplers are significantly more effective in reducing the buildup of longer-ranged correlations with $T$ than transmon couplers (whose parameters are optimized to minimize the $ZZ$ interaction). To fully confirm that this is a general difference between the two coupling schemes, one would have to check additional frequency patterns, however.
Figure 6: (a) Approximate ZZZZ coupling $\zeta$ defined in Eq. (21) between qubits $3$ and 44 as a function of external flux and capacitive coupling strength. (b) ZZZZ coupling for the isolated qubit pair. The avoided level crossing responsible for the jump in $\zeta$ is shown in panel (c).
@@
where $E_{n_{1}n_{2}}^{\prime}$ is the energy of the state, in which the qubits connected to the coupler are in the states $n_{1}$ and $n_{2}$, and all other qubits are in the 0 state. This becomes equal to the absolute values of nearest-neighbor couplings in Eq. (2) when higher-order terms involving 3 or more qubits vanish. As an example, Fig. 6(a) displays $\zeta$ as a function of $\phi_{\text{ext}}$ and $T$ for the qubit pair $2$ and $3$. For small $T$, the difference compared to the isolated qubit dimer [Fig. 6(b)] is mostly a change of the optimal value for $\phi_{\text{ext}}$. Around $T\approx 10$ MHz, however, there is a sudden increase of $\zeta$ caused by an avoided crossing between the eigenstates for $\bm{n}=01100000$ and $\bm{n}=00110000$ [Fig. 6(c)], showing that the effect of spectator qubits cannot always be accounted for by a simple shift of $\phi_{\text{ext}}$. It should also be noted that the optimization only addresses the nearest-neighbor $ZZ$ interactions, so we can not reasonably expect it to remove all stray couplings for larger $T$, where longer-ranged and multi-qubit terms become important. Nevertheless, the DMRG simulations confirmed that minimizing $\zeta$ independently for each coupler while keeping the remaining model parameters fixed is sufficient to bring all interaction strengths $|w_{\bm{n}}|$ below 50 Hz for weak capacitive coupling $T\leq 8$ MHz.
@@ IV Fluxonium with transmon couplers
When balancing the parameters of fluxonium qubits, several aspects must be considered. A sufficiently small $E_{C}$ allows for larger coupling capacitances, while $E_{L}$ should not be chosen too large in order to avoid smaller coherence times. In this specific example, we employ a quasi-periodic arrangement of qubit frequencies to mitigate the risk of frequency crowding. We emphasize, however, that there exists a broad parameter regime, and the precise choice may ultimately depend on the requirements of a specific gate scheme. Based on these considerations we propose the set of parameters listed in Table 2. The strengths of the capacitive couplings are again uniform. For each $T$, we first determine the value of $T_{c}$ that minimizes the average $ZZ$ coupling between nearest-neighbor qubits when treating the qubit dimers as isolated (see Fig. 7). We then use the parameters in Table 2 along with the optimized coupling $T_{c}$ for the MPS simulation of the full system.
Figure 8 shows the obtained nearest-neighbor $ZZ$ couplings and the average IPR. The $ZZ$ couplings are on the order of a few kHz and deviate only slightly from the results for the isolated qubit dimers. Moreover, the additional longer-ranged interactions and those involving three or more qubits have coefficients $w_{\bm{n}}$ with magnitude below 10 Hz, and are thus negligible. The IPR remains at a large value >0.9>0.9 for all considered parameters, indicating that the computational basis states remain strongly localized. This can likely be attributed to the fact that the capacitive coupling only weakly mixes the $|0\rangle$ and $|1\rangle$ states of the fluxonium, and that the additional transmons have a much larger transition frequency, so that the hybridization with the couplers is also suppressed.
@@
[…]
Figure 7: ZZZZ coupling averaged over decoupled subsystems of nearest-neighbor fluxonium qubits and the connecting transmons as a function of the coupling strengths $T$ and $T_{c}$. The line shows the $T$-$T_{c}$ curve used for the MPS simulations of the full system.
@@
V Discussion
Building a superconducting quantum computer requires achieving fast gate operations while minimizing unwanted crosstalk across the qubit chip. Although increasing coupling strengths enables faster gates, it also amplifies crosstalk and may even induce chaotic instabilities in larger systems, making it challenging to strike an optimal balance. For fluxonium qubits, we showed that a transmon-based tunable coupler can suppress $ZZ$ crosstalk to the level of $\sim 1\,$kHz. However, to contextualize this result, particularly in comparison to transmon architectures, we must assess the coupling strengths required for implementing gates with similar speeds. In Ref. [14], a coupling strength of $T\approx 120\,$MHz was used to implement a 60 ns CZ gate for two Fluxoniums. In contrast, a similar CZ gate in a transmon system with transmon tunable couplers, as demonstrated in Ref. [9], was achieved in a 38 ns gate time with $T\approx 20\,$MHz. Notice that both gates were studied in isolated two-qubit systems. Our simulations suggest that in a scaled-up qubit layout, residual crosstalk in a fluxonium system would be an order of magnitude lower, indicating that static $ZZ$-crosstalk might not be the primary bottleneck for scaling fluxonium-based architectures. For a more comprehensive study of the scalability perspectives of fluxonium architectures, it is of course important to also investigate other sources of cross-talk, e.g. between the dc flux lines for controlling the qubits and couplers. It should also be noted that by treating the nearest-neighbor coupling as an independent parameter, we did not account for capacitive loading [39], which can become a problem when scaling up the number of qubits as it limits the attainable coupling strengths for fluxoniums with higher connectivity.
In our simulations, we neglected the longer-ranged terms in the Hamiltonian that should become more relevant in both transmon and fluxonium systems as the coupling capacitances increase. As an example, consider the transmon architecture with tunable couplers at capacitive coupling strength T=10T=10\,MHz. Inverting the capacitance matrix that (approximately) corresponds to this model yields next-nearest-neighbor transverse couplings below 100100\,kHz. In perturbation theory, this causes corrections to the next-nearest-neighbor $ZZ$ interactions $w_{\bm{n}}$ in the range 10 Hz–100 Hz, which is on the order of those coefficients $w_{\bm{n}}$ in the model with only nearest-neighbor couplings, but well below the nearest-neighbor $ZZ$ interactions. The longer-ranged couplings therefore do not appear to significantly affect our results. However, because of frequency collisions, perturbation theory might underestimate the corrections for some qubit pairs. Here, further numerical analysis would yield more accurate quantitative predictions for these corrections.
Lastly, let us comment on the geometry of the system. Although the $2\times 4$-ladder considered in this work is small compared to an actual large-scale device, it already has an important feature of two-dimensional systems in that it includes loops with multiple coupling paths between qubits. In the regime we are interested in, where the qubits are localized but first signs of delocalization appear, we therefore expect the ladder model to give meaningful results for the $ZZ$ crosstalk in qubit lattices. The larger connectivity in a two-dimensional system may affect localization measures like the IPR, but since we always used the same geometry, the comparisons between architectures should still be valid.
VI Summary
We used density-matrix renormalization-group simulations to investigate how well different schemes for superconducting qubit devices are able to suppress the static $ZZ$ interaction when the system is scaled up. In agreement with previous works, we found that lattices of fixed-frequency transmons are limited to small capacitive couplings on the order of a few MHz, before the nearest-neighbor $ZZ$ interaction becomes prohibitively large. Although the addition of tunable couplers can reduce the effective interactions between qubits significantly, the qubits will still delocalize at sufficiently large capacitive couplings, and we indeed observe emerging long-range interactions due to near-resonant next-nearest neighbor qubits. Avoiding such frequency collisions should therefore also be beneficial for architectures with tunable couplers.
@@ VI Summary
For the architecture with fluxonium qubits and tunable transmon couplers, the results are promising, showing an efficient suppression of the $ZZ$ interaction and a large IPR over a wide range of capacitive coupling strengths.
As direction for future study, it would be interesting to investigate how longer-ranged couplings influence the delocation as coupling strengths are increased. In principle, such terms could be easily included in the DMRG simulations. It would also be possible to extend the MPS approach used here to time-dependent simulations [40], e.g., to investigate the effect of stray couplings during gate operations [41].
@@ Acknowledgements.
Data availability statement
All data that support the findings of this study are included within the article.